In 1980, Denmark had a GDP of $70 billion (measured in U.S. dollars) and a population of 5.1 million. In 2000, Denmark had a GDP of $160 billion (measured in U.S. dollars) and a population of 5.3 million. By what percentage did Denmark’s GDP per capita rise between 1980 and 2000?
1
Expert's answer
2020-11-17T07:17:18-0500

GDP per capita "= \frac{GDP}{Population}"

Denmark had a population of 5.1 million and GDP of $70 billion in 1980.

GDP per capita of Denmark in 1980 "= \frac{70 \;billion}{5.1 \;million} = 13725.49"

Denmark had a population of 5.3 million and GDP of $160 billion in 2000

GDP per capita of Denmark in 2000 "= \frac{160 \;billion}{5.3 \;million} = 30188.68"

Percentage change in GDP per capita between 1980 and 2000 is given below:

Percentage change "= \frac{GDP \;per \;capita \;of\; Denmark \;in \;2000 - GDP \;per \;capita \;of \;Denmark \;in \;1980}{GDP \;per \;capita \;of \;Denmark\; in \;1980} \times 100"

"= \frac{30188.68 – 13725.49}{13725.49}\times 100 = 120" %

GDP of Denmark was increased by 120 % between 1980 and 2000.